OpenStudy (anonymous):

It takes a person 30 minutes to gather 1 basket of fruit. Working at the same rate, how many baskets can 25 people gather in 8 hours?

OpenStudy (anonymous):

16 times 2 times 25

OpenStudy (anonymous):

sorry t16 times 1 times 25

OpenStudy (anonymous):

400

OpenStudy (anonymous):

firstly you have to calculate how much time 25 people will take to fill a single bucket. 1 person take = 30 25 person will take = 30/25 now we have 8 hours menas 8*60 = 480 min. so, the buckts that can be fill in this time will be = 480/(30/25) that is 480*25 /30 = 400